OpenStudy (anonymous):

The length of a rectangle in inches is sixteen less than seven time the width. If the perimeter of the retangle is 48 inches, what is the lengeh, in inches? a. 5 b. 8 c. 19 d. 40

OpenStudy (anonymous):

a=5

OpenStudy (anonymous):

L= 7w-16 L+L+W+W = 48 7w-16 + 7w-16 +w+w=48 12w -32 = 48 12w=80 w=5